OpenStudy (anonymous):

PLS pls pls help The polynomial P(x) is divided by (x-1)(x+2). Find the remainder if, given that P(1)=2 and P(-2)= 5

OpenStudy (anonymous):

@satellite73 help pls pls pls. I been stuck on this one for days...

OpenStudy (amistre64):

do you recall the remainder thrm?

OpenStudy (anonymous):

yea P(x) = (x-1)(x+2) Q(x) + R

OpenStudy (amistre64):

Since the P(1) and P(-2) are not zero, we know that the given is not a divisor right?

OpenStudy (anonymous):

ye

OpenStudy (amistre64):

what ideas have you run thru so far?

OpenStudy (anonymous):

i tried to sub P(1) and P(-2) into (x-1)(x+2) Q(x) + R

OpenStudy (amistre64):

yeah, all that really tells us is that R = 1 or -2 :)

OpenStudy (amistre64):

we are given 2 points, the least poly to fit between 2 points is a line, right?

OpenStudy (anonymous):

oh yea...

OpenStudy (amistre64):

and since the given is an x^2, the top would have to be the remainder ... of at most x

OpenStudy (amistre64):

you get where this is going?

OpenStudy (anonymous):

kinda...

OpenStudy (amistre64):

\[\frac{P(x)}{(x-1)(x+2)}=Q(x)+\frac{R(x)}{(x-1)(x+2)}\]

OpenStudy (amistre64):

R(x) is a poly of at most, degree 1; and since its value changes, it cant be a constant and must be a line

OpenStudy (amistre64):

the line between 1,2 and -2,5

OpenStudy (anonymous):

oh....

OpenStudy (amistre64):

im assuming you know how to define that line ...

OpenStudy (anonymous):

with gradient -1 right?

OpenStudy (amistre64):

maybe, i hadnt quite gotten to the line :)

OpenStudy (anonymous):

then i got line 3-x

OpenStudy (amistre64):

thats a great line

OpenStudy (anonymous):

would that really be the answer?

OpenStudy (amistre64):

and any multiple of it would work as well

OpenStudy (amistre64):

\[R(x)=k(3-x)\]

OpenStudy (amistre64):

the k migh tbe a little much tho; 3-x is sufficient

OpenStudy (anonymous):

yea thanks so much!!!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

you're* LOLOLOL

OpenStudy (amistre64):

.... they dont put that key in an opportune spot ;)